byeeee:

Which of the following exponential functions goes through the points (1, 12) and (2, 36)? f(x) = 4(3)−x f(x) = 3(4)−x f(x) = 3(4)x f(x) = 4(3)x

dude:

Right off the bat, lets look at the points. Do they go up or down?

byeeee:

up

dude:

So we can eliminate \(f(x) = 4(3)^{−x}\) and \(f(x) = 3(4)^{−x}\)

byeeee:

yeah

dude:

To check which one is right we can substitute each point into the equations (If you dont have access to a graphing calculator) Ex. (To check for answer choice D) \(f(x)=4(3)^x=>~~12=4(3)^1\) and \(36=4(3)^2\) If you do it should look something like this https://www.desmos.com/calculator/xvcyxstsha

byeeee:

ohh okay

byeeee:

hm

byeeee:

now what?

dude:

Do can you use graphing calculators?

byeeee:

no i can't

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude To check which one is right we can substitute each point into the equations (If you dont have access to a graphing calculator) Ex. (To check for answer choice D) \(f(x)=4(3)^x=>~~12=4(3)^1\) and \(36=4(3)^2\) \(\color{#0cbb34}{\text{End of Quote}}\) Then let's solve for these \(12=4(3)^1\) \(36=4(3)^2\)

byeeee:

im sorry i have to go now thanks for everything you've helped with ill be on later bye!

dude:

Alright

byeeee:

@dude hey could you help?

byeeee:

nvm i think my answer is D thanks for everything though!